OpenStudy (anonymous):

Finding Probability Using a Normal Distribution The weights of newborn baby boys born at a local hospital are believed to have a normal distribution with a mean weight of 3537 grams and a variance of 285,156. If a newborn baby boy born at the local hospital is randomly selected, find the probability that the weight will be less than 4172 grams. Round your answer to four decimal places.

OpenStudy (anonymous):

Been a while since I took stat, but I think this is a z-score problem

OpenStudy (anonymous):

Yes it is i need help desperately

OpenStudy (lyrae):

Fundamental probability theory states: \[P_{\chi}(\chi \le x) = F_{\chi}(x)\]In this question \(\chi\) is distributed as a normal distribution with \(\mu = 3537\) and \(\sigma^2 = 285,156\). \(F_{\chi}(x)\) is the cumulative distribution function of \(\chi\), i.e. normalcdf.

OpenStudy (anonymous):

So how do i do it?

OpenStudy (anonymous):

So how do i do it?

OpenStudy (lyrae):

Use a statistical calculator, Excel, wolfram alpha, etc. and use the cumulative distribution function for normal distribution (normalcdf) with given values for parameters x, \(\mu\) \(\sigma\).

OpenStudy (anonymous):

I dont understand what u just said at all

OpenStudy (anonymous):

I dont understand what u just said at all

OpenStudy (anonymous):

How do i find the standard deviation?

OpenStudy (lyrae):

Use this: http://www.danielsoper.com/statcalc3/calc.aspx?id=53 Mean = \(\mu\) = 3537 Standard deviation = \(\sqrt{variance}\) = \(\sqrt{\sigma^2}\) = \(\sigma\) = \(\sqrt{285,156}\) x = 4172

OpenStudy (anonymous):

Im sorry i still dont understand. I dont get that calculator. There isnt a a place for me to put the variance in

OpenStudy (anonymous):

or is the variance the standard deviation?

OpenStudy (lyrae):

standard deviation = \(\sqrt{variance}\)